OpenStudy (anonymous):

Last few questions and I can finally get some sleep! <3 1. How much heat energy is absorbed by 5.1 x 10^2 g of ice at -20°C as it is converted to water at 0°C? 2. What mass of water at 100°C can be converted to vapor at 100°C by the addition of 5.0 x 10^4 J of energy? 3. If 1.04 grams of NH3 are reacted, how many molecules of H2O are produced? 4NH3 + 7 O2 —> 4 NO2 + 6 H2O 4. When a 28.7 gram sample of KI dissolves in 60.0 grams of water in a calorimeter, the temperature drops from 27.2°C to 13.2°C. Calculate the Detla H for the process. KI (s) —> K+ (aq) + I- (aq)

OpenStudy (anonymous):

@Abhisar & @Somy Feel free to help if you can :p Most of this is going over my head as I missed some much class. Thanks! :)

OpenStudy (somy):

q 1 use this formula \[q = mc \Delta T\]

OpenStudy (somy):

c is constant which is equal to 4.18

OpenStudy (somy):

\[\Delta T\] is change in temperature as for the q1 its changing from -20 to 0 so finial- initial = 0- (-20) = +20 make sure your signs right because they show if the energy is released or absorbed

OpenStudy (somy):

im getting 42636 J

OpenStudy (somy):

same formula used in q 2 but now u do know q value which is Joules but u do not know mass so make mass as subject

OpenStudy (somy):

im getting 119.6g for this question 2

OpenStudy (somy):

in q 3 first find mole of NH3 u are given mass and Mr you can find from periodic table

OpenStudy (eric_d):

It shld be 0.06mol

OpenStudy (somy):

consider coefficient of NH3 which is 4

OpenStudy (eric_d):

so, 0.06/4= mol

OpenStudy (somy):

\[1.04 \div (4\times17) = 0.015 moles of NH3 \]

OpenStudy (somy):

ration between NH3 and H2O is 4:6 \[4 \rightarrow 0.015 moles\] \[6 \rightarrow X moles \] so mole of H2O will be \[(6 \times 0.015) \div 4 = 0.0225 moles\]

OpenStudy (somy):

if your q in 3 is how many molecules of H2O are produced then multiply mole of H2O by Avogadro's number which is \[6.02\times 10^{23}\]

OpenStudy (anonymous):

Wow thanks.. Sorry I fell asleep. @Somy

OpenStudy (somy):

it's ok i figured as much :)

OpenStudy (anonymous):

so for question 3 I do what?

OpenStudy (somy):

find mole of NH3

OpenStudy (anonymous):

How do I do that?

OpenStudy (somy):

mole= mass/Mr

OpenStudy (anonymous):

whats mr? Do I need the periodic table for this?

OpenStudy (somy):

yes

OpenStudy (somy):

Mr is molecular mass

OpenStudy (anonymous):

ohhhh ok. so I need to add up the mr from the periodic table, what is the other mass from?

OpenStudy (somy):

it's given to u in the question

OpenStudy (anonymous):

Ok so 17.031g?

OpenStudy (somy):

?? u have grams given in the question

OpenStudy (somy):

look in the question the mass given for NH3 is 1.04g u need to find Mr and then plug in the values into the formula :)

OpenStudy (anonymous):

17.031 g/mol is the molecular mass? right?

OpenStudy (anonymous):

So I got .061 moles NH3?

OpenStudy (somy):

Mr of N is 14 and of H is 1 u have 1 N and 3 H so 14+3= 17 u have coefficient 4 before NH3 in the reaction so 4*17= 68

OpenStudy (somy):

oh yeah 17.031 is also right, i just rounded of :)

OpenStudy (somy):

no it's gonna be 0.015moles

OpenStudy (anonymous):

o.O Lost me again. I suck at chemistry. How'd you get .015?

OpenStudy (somy):

look we said that coefficient of NH3 is 4 that means 4*17 = 68 -this is ur total Mr now mole =mass/Mr =1.04/68=

OpenStudy (somy):

4NH3 + 7 O2 —> 4 NO2 + 6 H2O here in reaction, u see 4 before NH3 right?

OpenStudy (anonymous):

yeah, makes sense. So is .015 the final answer?

OpenStudy (somy):

yes thats for mole now see the ratio between NH3 and H2O

OpenStudy (somy):

look at coefficients of each

OpenStudy (anonymous):

wait, so how many molecules of H2O are produced?

OpenStudy (somy):

u have to find moles of H2O first

OpenStudy (anonymous):

By the way - Sorry to be so quick and jumpy. I'm trying to finish this and get out the door so I'm not late.

OpenStudy (somy):

its ok look ration between NH3 and H2O is 4:6 (look at the equation u'll get it) so if 4 give 0.015moles how much will 6 give?

OpenStudy (somy):

here u do normal cross multiplication as in math

OpenStudy (anonymous):

.0225?

OpenStudy (somy):

thats right

OpenStudy (somy):

now they are asking number of Molecules so multiply this mole by Avogadro constant 6.02*10^23

OpenStudy (anonymous):

So: .0225 x 6.02 x 10^23? That game me 1.3545e22

OpenStudy (somy):

thats right

OpenStudy (somy):

1.3545*10^22

OpenStudy (anonymous):

so that's the final answer?

OpenStudy (somy):

yes

OpenStudy (anonymous):

YAY! Thanks so much @Somy :) Do you know the answer to number 4 by chance? I have to go so I can't solve it.

OpenStudy (somy):

ah no i didn't solve it yet

OpenStudy (somy):

q= mcT q= 60* 4.18 * 27.2°C= 6821.76 J q= 60* 4.18* 13.2°C.= 3310.56 J notice temperature is decreasing so its exothermic reaction 6821.76-3310.56= 3511.2 J but since its exothermic then it'll be -3511.2 J

OpenStudy (somy):

you are welcome btw :)

OpenStudy (anonymous):

So delta H = -3511.2 J

OpenStudy (somy):

yes

OpenStudy (anonymous):

YAAAAAY! Thank you <3 I love you soooo much right now. You just saved me :) I have to go but lets keep in touch...

OpenStudy (somy):

sure lol :D no problems :) are the answers in options though?

OpenStudy (anonymous):

No, I wish!

OpenStudy (somy):

oh that's sad, u didn't give the options so i did the questions the way i thought of them

OpenStudy (somy):

im sorry though T_T mb i made mistakes

OpenStudy (anonymous):

wait, made mistakes on what?

OpenStudy (somy):

no, i mean if none of my answers are matching your options

OpenStudy (somy):

mb i understood the questions wrong or made mistakes ( though i don't really see mistakes )

OpenStudy (anonymous):

@Somy Oh no! I was saying these questions weren't multiple choice. There weren't any options. Just had to solve it.

OpenStudy (somy):

oh lol good then :) i thought i did all wrong for u T_T